OpenStudy (anonymous):

Medal & Fan reward if answered correctly! Simplify completely the quantity 4 times x to the third power plus 5 times x to the 2nd power plus 3 times x all over x.

OpenStudy (anonymous):

use equation editor here || || \/

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

you see the blue button that says "\(\Sigma\) Equation "?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

really? looks right below the box you just typed in

OpenStudy (anonymous):

that was you can make it look like \[\frac{4x^3+5x^2+2x}{x}\]

OpenStudy (anonymous):

oh its a timed test and i really need help.

OpenStudy (anonymous):

3x3 + 5x + 2 3x2 + 5x2 + 2x 3x5 + 5x3 + 2x2

OpenStudy (anonymous):

reduce each exponent by 1

OpenStudy (anonymous):

Those are the possible answers^

OpenStudy (anonymous):

go with the middle one

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (anonymous):

oh no! go with the top one

OpenStudy (anonymous):

Wait lol, i gave you the wrong possible outcomes.

OpenStudy (anonymous):

\[\frac{3x^3+5x^2+2x}{x}\]\[=\frac{3x^3}{x}+\frac{5x^2}{x}+\frac{2x}{x}=3x^2+5x+2\]

OpenStudy (anonymous):

4x2 + 5x + 3 the quantity 4 times x to the 2nd power plus 5 times x plus 3 all over x 4x2 + 5x2 + 3x 4x4 + 5x3 + 3x2

OpenStudy (anonymous):

ok then \[4x^2+5x+3\] i thought the answers looked strange

OpenStudy (anonymous):

ok now i am lost

OpenStudy (anonymous):

Why are you lost?

OpenStudy (anonymous):

this is what your picture (gif) was \[\frac{4x^3+5x^2+2x}{x}\]

OpenStudy (anonymous):

is it a different question, or this one?

OpenStudy (anonymous):

4x2 + 5x + 3 the quantity 4 times x to the 2nd power plus 5 times x plus 3 all over x 4x2 + 5x2 + 3x 4x4 + 5x3 + 3x2

OpenStudy (anonymous):

The question & the possible answers

OpenStudy (anonymous):

then the answer is \[4x^2+5x+3\]

OpenStudy (anonymous):

Okay I have three more, could you possibly help me with them?